Verify rolle's theorem

f(x)=(x-a)^m(x-b)^n,where m,n are positive intigers, in the interval [a,b]


1
Expert's answer
2021-07-12T09:28:59-0400

Rolle's theorem state that any real value differentiable function which gives equal value at two distinct points will have a point in between them where "f^{'}(x) =0" exist .


Now given function is "f(x)=(x-a)^{m}(x-b)^{n}" , where m , n are positive integers , in the interval [a,b].


Now putting "x=a," we get "f(a)=0"


putting "x=b" , we get "f(b)=0" , so hence it proves that rolle's theorem satisfy .


Now , we have to check "f^{'}(x)" is 0 between the point a and b respectively .


Now , in this case we will two functions the first one will be "(x-a)^{m}" and the second one will be "(x-b)^{n}."


"f^{'}(x)=\dfrac{d[(x-a)^{m}(x-b)^{n}]}{dx}=(x-b)^{n}m(x-a)^{m-1}+(x-a)^{m}n(x-b)^{n-1}"



now putting "f^{'}(x)=0" , we get ,


"=m(x-a)^{m-1}(x-b)^{n}+(x-a)^{m}n(x-b)^{n-1}=0"


"=(x-a)^{m-1}(x-b)^{n-1}[m(x-b)+n(x-a)]=0"


"(m+n)x=mb+na"


"x=\dfrac{(mb+na)}{m+n}"


"=\dfrac{(mb+na)}{m+n}>a" , m and n are given that are positive numbers .


"=mb+na>ma+na\implies" "m(b-a)>0"


Similarly ,


"=\dfrac{(mb+na)}{m+n}<b"


"=mb+na<mb+nb\implies n(a-b)<0"


So it implies that


"a<\dfrac{(mb+na)}{m+n}<b"


So , hence our rolle's theorem verifies.